Write an expression with a variable for each situation described below. Then write a phrase telling what the variable represents.
1. Jason had 12 cookies and then ate some.
2. Loretta had 5 books and then bought some more.
3. Hal had 8 pens. He bought 5 more and then gave away some.

Write an expression, then evaluate the expression.
4. Barbara saved d dollars a week from her job for 6 weeks. How much did Barbara save? Evaluate the expression for d = $18.75.
5. Tom had $40 before he went shopping. After he bought some school supplies, he had n left. How much did Tom spend on school supplies? Evaluate the expression for n = $15.

Answer by rfer(16322)

1)
12-x=
2)
5+x=
3)
8+5-x=
4)
6d=
6*18.75=$112.50
5)
40-n=
40-15=$25